A sphere and a cube have the same volume.If the radius of the sphere is 5.00 cm,then what is the surface area of the sphere?
1
Expert's answer
2020-09-28T08:09:04-0400

The surface area of the sphere

"A=4\pi R^2=4\times 3.14\times (5.00\:\rm cm)^2=314\: \rm cm^2"